# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### v27

@@ 7.3 Information Efficiency of Prices
Proposition 7.12 is intuitive and consistent with the comparative statics of $\alpha^{*}$. As the market maker’s prior becomes more dispersed over possible signals, he is less able to infer information from trades, which lowers the information efficiency of prices. The marginal effect is diminishing because the larger trades the insider makes to exploit the additional prior uncertainty of the market maker also increases the variation of his trades across signals. This increased variation reveals some of his information, partially offsetting the loss in information efficiency.
@@ 8 Empirical Implications
Computation: For an order flow realization $\omega$, we use the pricing kernel $P^{*}(,\cdot,|\omega;s_{1})$ to compute the equilibrium option prices across strike $K$. Based on these option prices, we compute the implied volatility $\sigma_{\scriptscriptstyle IV}(\omega,K)$ for each strike $K$. This is repeated for $1,000$ realizations of $\omega$, with the plots displaying the average implied volatilities over log-moneyness.
The risk-free rate and dividend yield are set to zero.
@@ A.1 Model Assumptions
The marginal utility functional $\partial\!\kern 0.80002ptJ(\,\cdot\,;W)\colon\mathbb{R}^{\infty}\rightarrow\mathbb{R}$ is the Gâteaux-derivative of the expected utility functional $J(\,\cdot\,|s_{k};\widetilde{W})\colon\mathbb{R}^{\infty}\rightarrow{\mathbb{R}}$ at $W$. (Like $J(\,\cdot\,|s_{k};\widetilde{W})$, its Gâteaux-derivative $\partial\!\kern 0.80002ptJ(\,\cdot\,;W)$ is understood to be conditional on $s_{k}$ and parameterized by $\widetilde{W}$.) We need to show this Gâteaux-derivative decomposes into
@@
Because the Gâteaux derivative of the payoff functional
$W\,\mapsto\,\sum\limits_{i}W(x_{i})\eta(x_{i}|s_{k}),\,\mathbb{R}^{\infty}\rightarrow\mathbb{R},$
is trivially identified with $\eta(\,\cdot\,|s_{k})$, the characterization of the marginal payoff functional $\partial\!\kern 0.80002ptJ_{p}(\,\cdot\,;W)$ in Equation (15) is immediate.
Now, consider the cost functional
$J_{c}(W)=\sum_{i}W(x_{i})\overline{P}(x_{i},W;\widetilde{W}).$
To make the dependence of expected AD price $\overline{P}(x_{i},W;\widetilde{W})$ (defined as in Equation (14)) on $W$ more explicit, we write out

### v28

@@ 7.3 Information Efficiency of Prices
Proposition 7.12 is intuitive and consistent with the comparative statics of $\alpha^{*}$ shown in Figure 2. A larger number of signals makes it more difficult for the market maker to infer information from trades, thereby reducing the information efficiency of prices. The marginal effect on information efficiency is diminishing because the resulting larger trades by the insider, aimed at exploiting the market maker’s prior uncertainty, increase the variation in his trades across signals. This increased variation reveals some of his information, partially offsetting the loss in information efficiency.
@@ 8 Empirical Implications
Computation: For an order flow realization $\omega$, we use the pricing kernel $P^{*}(,\cdot,|\omega;s_{1})$ to compute the equilibrium option prices across strike $K$. Based on these option prices, we compute the implied volatility $\sigma_{\scriptscriptstyle IV}(\omega,K)$ for each strike $K$. This is repeated for $1,000$ realizations of $\omega$, with the plots displaying the average implied volatilities over log-moneyness. The risk-free rate and dividend yield are set to zero.
@@ A.1 Model Assumptions
The marginal utility functional $\partial\!\kern 0.80002ptJ(\,\cdot\,;W)\colon\mathbb{R}^{\infty}\rightarrow\mathbb{R}$ is the Gâteaux-derivative of the expected utility functional $J(\,\cdot\,|s_{k};\widetilde{W})\colon\mathbb{R}^{\infty}\rightarrow{\mathbb{R}}$, as defined in Equation (14), at $W$. (Both $J(\,\cdot\,|s_{k};\widetilde{W})$ and its Gâteaux-derivative $\partial\!\kern 0.80002ptJ(\,\cdot\,;W)$ are conditional on $s_{k}$ and parameterized by $\widetilde{W}$.) $J(W|s_{k};\widetilde{W})=J_{p}(W)-J_{c}(W)$ is the difference between
[…]
and
[…]
We need to show its Gâteaux-derivative decomposes into
@@
Because the Gâteaux derivative of the payoff functional $J_{p}(\,\cdot\,)$ is trivially identified with $\eta(\,\cdot\,|s_{k})$, the characterization of the marginal payoff functional $\partial\!\kern 0.80002ptJ_{p}(\,\cdot\,;W)$ in Equation (15) is immediate.
Now, consider the cost functional $J_{c}(\,\cdot\,)$. To make the dependence of expected AD price $\overline{P}(x_{i},W;\widetilde{W})$ on $W$ more explicit, we write out